OpenStudy (anonymous):

can someone explain to me what this means? give a careful definition of the trigonometric functions defined on "proper angles"(angles which can be part of a right triangle)

OpenStudy (anonymous):

I think I might be able to. Lemme just double check the definition of a proper angle to be on the safe side.

OpenStudy (anonymous):

whoops wrong question lol

OpenStudy (anonymous):

Hmm... Don't see a definition for that. But... Sine of an angle, let's call it "theta", will be the number you get by dividing the side that is on the opposite side of the triangle by the hypotenuse (the longest side, it's opposite of the 90 degree angle). So, if your triangle is a right triangle and has these sides: 3,4,5 and you pick the angle that is opposite of the side that has 3, your sin of that angle would be 3/5. The cosine of the angle is the adjacent side (the touching side that is not the hypotenuse) divided by the hypotenuse. So for our triangle it is 4/5. Finally, the tangent will be the only remaining combo: opposite and adjacent. The tangent of an angle is the answer that you get when you divide the opposite by the adjacent, or 3/4 for our triangle. Hope this is what you wanted?

OpenStudy (anonymous):

hey liz u got it?

OpenStudy (anonymous):

kind of. so sin(t)=y, cos(t)=x, tan(t)=y/x where x cannot=0. So sin of an angle equals opposite over hypotenuse=y/1=y and cosine of an angle= adjacent over hypotenuse=x/1=x and tangent of an angle=opposite over hypotenuse=y/x where x cannot =0. Is this the best way to explain it using these words

OpenStudy (anonymous):

sine is opposite by hypotenuse cos is adjacent by hypotenuse tan is opposite by adjacent

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

I do a little. I have in my notes. Any ratio we make from the side lengths of a right triangle is invariant under similarity, so depends on the angles, not the actual triangle. what does this mean

OpenStudy (anonymous):

it means the ratio is same for a given angle anywhere

OpenStudy (anonymous):

That works, dizliz. The reason you can use x or y is because you can draw a triangle on a coordinate plane (or a graph as it's more commonly called). But it's a good idea to remember opposite/hypotenus/adjacent as that's how triangles are more commonly encountered. A trick to remember them: cos without the o is cs. cs sounds like x. So it's x/hypotenuse. sin sounds like sYn. So sine is Y/hypotenuse. :)

OpenStudy (anonymous):

does that mean that if we enlarge the triangle that the ratio is the same

OpenStudy (anonymous):

yes surely

OpenStudy (anonymous):

That's right!

OpenStudy (anonymous):

so which angles can be part of a right triangle

OpenStudy (anonymous):

any angle can be

OpenStudy (anonymous):

Thing of it like this: No matter how big you make a square, they will always have the same ratio to the sides (as long as the angles remain the same). They'll always be 1:1:1:1. Likewise, if you have an equilateral triangle, the sides will always be 1:1:1. If you have a 30:60:90 triangle, the sides will always be 1:2:3 (divide each amount by 30).

OpenStudy (anonymous):

Any angle below 90 can be.

OpenStudy (anonymous):

30, 60, 90 triangle doesnt have sides in 1:2:3

OpenStudy (anonymous):

@Him: If the angle is 90 or more, then it's no longer a real/proper triangle.

OpenStudy (anonymous):

Oh right. It's 1:2:sqrt(3) for my example if I recall correctly.

OpenStudy (anonymous):

dont confuse her now

OpenStudy (anonymous):

liz come to chat

OpenStudy (anonymous):

Did you have any other questions, Liz?

OpenStudy (anonymous):

I have a ton of questions. I have an exam tom at 9 and I have about four more qeustions on this take home quiz

OpenStudy (anonymous):

dont worry well help

OpenStudy (anonymous):

That is helpful, so I can draw a triangle to prove this

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

So if I draw a triangle with sides 1,2,3 and enlarge it by a factor of 2 and prove it works. Will that help answer this question

OpenStudy (anonymous):

Yeah, that'd work. Only difference would be that each side would be twice as long.

OpenStudy (anonymous):

The angles would still remain the same.

OpenStudy (anonymous):

yeah in ur first case u write down all the three ratios.....and then in the second case all ur sides are multiplied by 2..so the ratios wil remain the same..get it?

OpenStudy (anonymous):

do u understand liz?

OpenStudy (anonymous):

It's kind of like how no matter how big or small you make a square, they'll always be the same, as long as all the sides are enlarged by the same ratio. But if you decide to change the ratio for even one side, you'll either get a trapezoid or a rectangle or just a random quadrilateral.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

did u unerstand the half angle formula?

OpenStudy (anonymous):

No I need that explained to me. Give me one sec to finish this question

OpenStudy (anonymous):

u there?

OpenStudy (anonymous):

liz?

OpenStudy (anonymous):

Yes. Now how would i answer this question. Suppose f is any function defined on the set of all real numbers with the properties: f is one to one; f has image equal to the set of all positive real numbers; f satisfies the addition law f(x+y)=f(x) f(y) for all real numbers x and y

OpenStudy (anonymous):

whts the question?

OpenStudy (anonymous):

? dont take 12 weeks to reply please

OpenStudy (anonymous):

Sorry, show that f has an inverse function g, with domain the set of positive real numbers, and show that g satisfies g(xy)= g(x)+g(y) for all positive real numbers x and y

OpenStudy (anonymous):

my computer is acting funny again

OpenStudy (anonymous):

there are two ways to do this/....do u know derivatives?

OpenStudy (anonymous):

no. we used exponentials and logarithms

OpenStudy (anonymous):

tell me where you started off...cz the long part is obtaining the function from these conditions...did u go to exponents directly?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

so see the function f(x) = a^x is a function which satisfies ur conditions..check em one by one..and tell me doesnt it?

OpenStudy (anonymous):

funny comp agn?

OpenStudy (anonymous):

no just thinking

OpenStudy (anonymous):

I want to show f using exponents and g using its inverse log

OpenStudy (anonymous):

im coming to that..so are we decided that f satisfies ur conditions?

OpenStudy (anonymous):

Im not sure

OpenStudy (anonymous):

f(x) = a^x f(x)f(y) = a^(x) a^(y) = a(x+y) ok?

OpenStudy (anonymous):

k

OpenStudy (anonymous):

so now do u know how to find an inverse of a function?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so y = f(x) = a^x now find its inverse

OpenStudy (anonymous):

Not in this form. I know how to write log for 2^3=8

OpenStudy (anonymous):

see \[y = a^x , \ now \ take \ \log \ \to \ the \ base \ a\ on \ both \ sides\] you will get \[\log_{a}y = \log_{a} (a^x) => \log_{a}y = xlog_{a}a = x \]

OpenStudy (anonymous):

now interchange x and y so \[y = f(x) = \log_{a}x\]

OpenStudy (anonymous):

understand>

OpenStudy (anonymous):

??

OpenStudy (anonymous):

A little better

OpenStudy (anonymous):

dont worry now..tell me....next question....or any doubts?

OpenStudy (anonymous):

I feel like I can't make a connection with it

OpenStudy (anonymous):

I dont know what it is.

OpenStudy (anonymous):

functions and stuff?

OpenStudy (anonymous):

come to chat..its faster

OpenStudy (anonymous):

Could we try using something with a base 2 to explain from the beginning

OpenStudy (anonymous):

as you wish madam

OpenStudy (anonymous):

let \[y = f(x) = 2^x\] ok?

OpenStudy (anonymous):

come to chat

OpenStudy (anonymous):

Its not working

OpenStudy (anonymous):

it isnt

OpenStudy (anonymous):

so is the function ok?

OpenStudy (anonymous):

no my message isnt going through on chat

OpenStudy (anonymous):

neither is mine

OpenStudy (anonymous):

i need f and g to undo each other

OpenStudy (anonymous):

they will..if 2^x is f whts g? tell me?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

for instance I have 2^2=4 which is log2(x)=2. Is that correct

OpenStudy (anonymous):

yes..so the inverse fn is log2(x)..isnt it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so now see if f and g cancel out

OpenStudy (anonymous):

How do i do that

OpenStudy (anonymous):

see if g is the inverse of x then f(g(x)) = x...do u understand?>

OpenStudy (anonymous):

What numbers do I put in place of f, g,x

OpenStudy (anonymous):

f(x) is 2^x ....g(x) is log2(x) so f(g(x)) = 2^(log2 (x)) isnt it?

OpenStudy (anonymous):

which is equal to x

OpenStudy (anonymous):

almost there thinking about it

OpenStudy (anonymous):

ah, because 2 and log2 cancel each other out. But why is that true

OpenStudy (anonymous):

its a rule of logarithm

OpenStudy (anonymous):

how can i show that g(xy)=g(x) +g(y)

OpenStudy (anonymous):

bcoz log(xy) = log x + log y

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

no, how does that work. is it because the inverse exponentials with addition is multiplication

OpenStudy (anonymous):

yes